OpenStudy (anonymous):

i need help understanding and figuring this out -x+2y=2 graph the linear equation and it needs three solutions

OpenStudy (marissalovescats):

Put it into y=mx +b

OpenStudy (anonymous):

yeah graphing aint my thing

OpenStudy (marissalovescats):

I'll let someone else help you I don't think I'm quite on the right track haha

OpenStudy (anonymous):

but i have to plot and have 3 different solutions to slove the same linearr equation

OpenStudy (marissalovescats):

You mean just find 3 points on the graph?

OpenStudy (anonymous):

yeah and they all have to solve the equation

OpenStudy (marissalovescats):

Well if the points are on the line, they should be solutions.

OpenStudy (anonymous):

but their isnt any on the line thats wat i hve to figure out

OpenStudy (marissalovescats):

Are you given the solutions?

OpenStudy (anonymous):

no im given the equation and i have to figure out the x and y

OpenStudy (anonymous):

and there has to be 3 soulutions

OpenStudy (marissalovescats):

What Im saying is.. Put this equation into slope intercept form y=mx+b and graph it then pick 3 points on your line that you graph and those should be solutions for your equation.

OpenStudy (marissalovescats):

@jim_thompson5910

OpenStudy (anonymous):

\[y=1/2x+1\]

OpenStudy (marissalovescats):

Okay now graph it

OpenStudy (marissalovescats):

1/2 is your slope and 1 in you y-int

OpenStudy (jdoe0001):

is a straight line so for that all you really need is 2 points to plot it try setting y = 0 that way you get the x-intercept, the y-intercept is already in the mx+b equation, which is b = 1

jimthompson5910 (jim_thompson5910):

if x = 0, then y is.. -x+2y=2 -0+2y=2 0 + 2y = 2 2y = 2 y = 2/2 y = 1 So we know the point (0,1) is on the line -x+2y=2 ------------------------ If x = 2, then y is -x+2y=2 -2+2y=2 2y = 2 + 2 2y = 4 y = 4/2 y = 2 giving us the point (2,2) which is also on the line -x+2y=2

OpenStudy (marissalovescats):

Or you could do it there way :P

jimthompson5910 (jim_thompson5910):

you keep this up for however many points you want next you plot each point finally you draw a straight line through every point to graph -x+2y=2

OpenStudy (anonymous):

so my third piont is 0,4 ??

jimthompson5910 (jim_thompson5910):

(0,4) is not on the line

OpenStudy (anonymous):

4,4??

jimthompson5910 (jim_thompson5910):

-x+2y=2 -4+2y=2 ... plug in x = 4 2y = 2+4 2y = 6 y = 6/2 y = 3 when x = 4, y = 3 the point (4,3) is on the line

OpenStudy (anonymous):

o okaii that makes a whole alota since :) thanks

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

can u help me with some more??

jimthompson5910 (jim_thompson5910):

sure, one more

OpenStudy (anonymous):

x-3y=-6

jimthompson5910 (jim_thompson5910):

when x = 0, y = ???

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

when x = 3, y = ???

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

when x = 6, y = ???

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

so we have these 3 points (0,2), (3,3), (6, 4)

OpenStudy (anonymous):

omg U MADE THAT SO EASY FOR MEH :) thanks:)

jimthompson5910 (jim_thompson5910):

I didn't do much, you did all the work lol

jimthompson5910 (jim_thompson5910):

but you're welcome

OpenStudy (anonymous):

but still u broke it down and helped meh understand it better soo thanks

jimthompson5910 (jim_thompson5910):

I'm glad I helped out and that you understand it all

OpenStudy (anonymous):

me too